OpenStudy (anonymous):

Identify whether the series is a convergent or divergent geometric series and find the sum, if possible. @perl @inkyvoyd

OpenStudy (inkyvoyd):

if the absolute value of hte common ratio r (the number under the xponent) is greater than or equal to 1, then the geometric series diverges.

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty}12(3/5)^{i-1}\]

OpenStudy (inkyvoyd):

the formula for infinite geometric series is \(\Huge \sum_{i=1}^{\infty}a(r)^{i-1}=\frac{a}{1-r}\)

OpenStudy (inkyvoyd):

this ONLY applies if the absolute value of r is less than or equal to 1.

OpenStudy (anonymous):

so its divergent

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (inkyvoyd):

no @woohoo r=3/5. 3/5=0.6, and .6 is less than 1

OpenStudy (anonymous):

congernt

OpenStudy (anonymous):

convergent @inkyvoyd

OpenStudy (inkyvoyd):

yes. so then, use the provided formula.

OpenStudy (anonymous):

is it 30 @inkyvoyd

OpenStudy (inkyvoyd):

Yes, I think so.

OpenStudy (anonymous):

You think so, im feeling kinda skeptical

OpenStudy (anonymous):

?

OpenStudy (inkyvoyd):

If I were you, I'd put down that answer. PS: you might want to check the other question. I think we should've used n=5 and not n=6

OpenStudy (agent0smith):

We did one like this earlier... did you already forget it? :P

OpenStudy (anonymous):

this onw is diffrent @agent0smith

OpenStudy (anonymous):

one

OpenStudy (agent0smith):

No, it's the same, at least deciding whether the series is convergent or not. It the same as the one where you had 15(4)^(i-1).

OpenStudy (anonymous):

its same in where its asking convergent or not but the problem in of itself is different. :)

OpenStudy (agent0smith):

Yes, but i mean you did not correctly identify whether it's convergent or not. Had you looked at the past question, you would have been able to.

OpenStudy (anonymous):

ohh now I understand you, I'm kinda slow whe it comes to Math

OpenStudy (agent0smith):

If you write notes, or at least look back over your past questions when you find a similar one, that'll help the information sink in.

OpenStudy (anonymous):

Thannks anyway for all your help! I really appreciate it! @agent0smith